OpenStudy (anonymous):

certain # of tick to give away. 5 people got tik. the 1'st got one third of tick. the 2nd got one fourth of the tick. and the 3rd got one fifth of tick. the 4th got eight tick, aand there were five tick left for the fifth person. find the total # of tickets given away ........so ????about w dwn

OpenStudy (sandra):

1/3t + 1/4t + 1/5t + 8 + 5 = t

OpenStudy (sandra):

can you solve from there?

OpenStudy (anonymous):

iam getting there but i think i want to put all t together right by add?

OpenStudy (sandra):

right - it's combining like terms.

OpenStudy (sandra):

for example, if I give you 5oranges and 10oranges, how many oranges do you have?

OpenStudy (anonymous):

15 yes,,, is just the way the put it together sounds difficult

OpenStudy (sandra):

right exactly. so now you have to add the fractions by finding the lowest common denominator of 3, 4, and 5

OpenStudy (sandra):

well, or any common denominator really

OpenStudy (anonymous):

awww kk. ur good..lol... math is not my subj can u tell?

OpenStudy (sandra):

1. 1/3t +1/4t + 1/5t + 8 + 5 = t 2. 1/3t + 1/4t + 1/5t +13 = t Now 3x4x5 = 60. so we want to get a 60 on the bottom of each fraction by multiplying by different variations of 1 (since anything/anything = 1)

OpenStudy (sandra):

so 3. 1/3t * 20/20 + 1/4t * 15/15 + 1/5t * 12/12 + 13 = t

OpenStudy (sandra):

4. 20/60t + 15/60t + 12/60t + 13 = t

OpenStudy (sandra):

5. (20+15+12)/60t + 13 = t

OpenStudy (sandra):

6. 47/60t + 13 = t

OpenStudy (sandra):

7. 47/60t - 47/60t + 13 = t - 47/60t

OpenStudy (sandra):

8. 13 = 13/60t

OpenStudy (sandra):

13/(13/60) = (13/60t)/(13/60)

OpenStudy (sandra):

t = 60*13/13

OpenStudy (sandra):

t = 60

OpenStudy (sandra):

so now put 60 in place of t back into the equation to check it

OpenStudy (sandra):

60*1/3 + 60*1/4 + 60*1/5 + 8 + 5 = 60 ?

OpenStudy (sandra):

20 + 15 + 12 + 13 = 60, 60=60

OpenStudy (sandra):

so we were right

OpenStudy (anonymous):

yes got the same answ thank u

OpenStudy (sandra):

no problem. good luck!

OpenStudy (anonymous):

process was def long but well explain

OpenStudy (sandra):

heh yeah sorry - these are tough at first. the toughest part is turning the word problem into an equation, but it comes easier with practice

OpenStudy (anonymous):

i bet iam inthe process of receiving my textbook so for noe examples and notes

OpenStudy (anonymous):

but i appreciate ur patient and time